# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v2

@@ IV.3 Solution of the random-mode model
$C_{\star}^{\phi}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star}.$ (36c)
These equations are equivalent those for an i.i.d. model (Eqs. 6971) with coupling strength
@@
The four-point function is (Appendix I.2)
@@
where we have simplified notation by suppressing frequency variables via the shorthand $C^{\phi}_{12}=C^{\phi}_{\star}(\omega_{1})C^{\phi}_{\star}(\omega_{2})$ and $S^{\phi}_{12}=S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})$. The same solution is obtained through a more involved, but intuitive, two-site cavity calculation (Appendix L).
This solution for the random-mode model is one of our main results that we now spend some time interpreting.
@@ IV.4 Interpreting the solution
Figure 4: Dimension of activity and collective timescales in the random-mode model. (a) Activity dimension $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ for various coupling strengths $g_{\text{eff}}$. Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions. Inset: Extended $g_{\text{eff}}=6$ case, showing convergence to i.i.d. coupling behavior with growing $\alpha$. (b) Normalized four-point function $\overline{\Psi^{\phi}(\tau,0)}\equiv\Psi^{\phi}(\tau,0)/\Psi^{\phi}(0,0)$ for various $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$. Inset: Theory curves for all $\alpha\text{PR}^{D}$, demonstrating relative invariance of collective timescales to the effective rank. Simulations use $N=5000$ neurons.
@@
Our analysis reveals two features of connectivity that characterize how it shapes activity, namely, the effective coupling strength $g_{\text{eff}}$ and the effective rank $\alpha\text{PR}^{D}$. The former characterizes local structure (the size of individual couplings) while the latter characterizes global structure (the number of large connectivity components). In the equations determining $C_{\star}^{\phi}(\tau)$ and $\Psi^{\phi}_{\star}(\bm{\tau})$ above, we see immediately that global structure in the connectivity is invisible at the single-neuron level but shapes collective activity. Specifically, the correlation function $C^{\phi}_{\star}(\tau)$ is fully determined by $g_{\text{eff}}$ and is insensitive to the effective rank of connectivity. This has important implications for neural data analysis. In particular, analyses of single-neuron, non-collective activity properties may miss signatures of structured connectivity that only become apparent when analyzing collective activity properties among simultaneously recorded neurons.
@@
What is the nature of the dependence of collective-activity properties, such as the dimension of activity, on the effective rank? To plot this relationship and validate the theory, we simulated networks and computed the participation ratio from the empirical equal-time covariance matrix of the activations. We used a component-strength spectrum $D_{a}\propto\exp(-\beta_{D}a/M)$ which yields $\text{PR}^{D}=\beta^{-1}_{D}\tanh(\beta_{D})$. To attain a desired effective rank $\alpha\text{PR}^{D}<1$, we set $\alpha=M/N=1$ and solved for $\beta$. For $\alpha\text{PR}^{D}>1$, we fixed $\text{PR}^{D}=1$ and increased $\alpha$. We adjusted $g_{\text{eff}}$ by rescaling $D_{a}$ uniformly. We find excellent theory-simulation agreement (Fig. 4). Fig. 4a illustrates how the dimension of activity, $\text{PR}^{\phi}$, varies with $\alpha\text{PR}^{D}$ for different values of $g_{\text{eff}}$. Consistent with i.i.d.-coupling networks, $\text{PR}^{\phi}$ increases monotonically with $g_{\text{eff}}$ [30]. It also increases monotonically with $\alpha\text{PR}^{D}$.
@@
Taking $\alpha\text{PR}^{D}\to\infty$ for fixed $g_{\text{eff}}$ in Eq. 38 recovers the formula for i.i.d. couplings (Eq. 72) of [30]. Indeed, in this limit, $\bm{J}$ approaches an i.i.d. matrix. Correspondingly, $\text{PR}^{\phi}$ approaches the dimension of a network with i.i.d. couplings (Fig. 4a inset). Comparing Eq. 72 to the corresponding i.i.d. formula ($1/\alpha\text{PR}^{D}=0$) and noting that single-neuron properties are unaffected by $\alpha\text{PR}^{D}$, one sees that any structure beyond i.i.d. connectivity ($1/\alpha\text{PR}^{D}$ finite) strictly decreases the dimension of activity.
@@
The dimension of activity depends on $\Psi^{\phi}_{\star}(\bm{\tau})$ at $\bm{\tau}=(0,0)$; we now examine the temporal profile of this function. Fig. 4b shows a normalized version of $\Psi^{\phi}(\tau,0)$, which is related to the correlation functions of leading principal components, for various values of $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$. The decay timescale of $\Psi^{\phi}_{\star}(\tau,0)$ decreases with increasing $g_{\text{eff}}$, approaching a limiting behavior at $g_{\text{eff}}\approx 10$. While $\alpha\text{PR}^{D}$ affects the overall scale of this function, and consequently the dimension of activity, it has little effect on the decay timescale, and thus little effect on the timescales of leading principal components (Fig. 4b, inset). Thus, the effective rank primarily influences the dimension of activity rather than the collective temporal structure.
@@
A remarkable feature of these results is that only two connectivity features, $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$, are sufficient to determine the two- and four-point functions. Due to the nonlinear, high-dimensional nature of the network, this reduction is unexpected. Indeed, this reduction breaks down when mode overlaps are introduced (Sec. VII), in which case the formula for the four-point function requires knowledge of the full joint distribution of component strengths and overlaps.
@@ V Limiting behavior of the random-mode model
Here we examine the behavior of the random-mode model in two limiting cases: as the system approaches the transition to chaos from above ($g_{\text{eff}}\to 1^{+}$), and in the limit of low effective rank ($\alpha\text{PR}^{D}\to 0^{+}$).
We first consider the limit $g_{\text{eff}}\to 1^{+}$ for arbitrary $\alpha\text{PR}^{D}$. The relevant expansion parameter is $g_{\text{eff}}-1\ll 1$. For networks with i.i.d. couplings with coupling strenght $g$, it is known that, as $g\to 1^{+}$, $\Psi^{a}_{\star,\text{i.i.d.}}(\bm{\tau})\sim 1/(g-1)$ and $\text{PR}^{a}_{\text{i.i.d.}}\sim(g-1)^{3}$ in this limit [30]. For the random-mode model, to leading order in $1/(g_{\text{eff}}-1)$,
@@
The difference between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$ is a higher-order correction, so we also have $\text{PR}^{a}=\text{PR}^{S}\text{PR}^{a}_{\text{i.i.d.}}$. That is, in the limits of both low effective rank and small effective coupling strength, the dimension of activity is equal to the dimension of the singular-value spectrum times the dimension of activity in an i.i.d. network.
@@
We verify this limiting behavior in simulations using $\alpha=1$ and a step function for the component strengths: $D_{a}=\text{const.}$ for $a/N\leq\text{PR}^{D}$ and $D_{a}=0$ otherwise. We varied $\text{PR}^{D}$, observing linear relationships with $\text{PR}^{\phi}$ and $\text{PR}^{x}$ in the $\text{PR}^{D}\to 0^{+}$ limit (Fig. 10a).
Focusing on $K^{\phi}(g_{\text{eff}})$, note that its expression (Eq. 43) resembles the formula for the dimension of activity in an i.i.d. network with coupling strength $g_{\text{eff}}$ (Eq. 72), but includes an additional factor $|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}$ in the integral. Numerical evaluation confirms that, as expected from Eq. 41, $K^{\phi}(g_{\text{eff}})\to\text{PR}^{\phi}_{\text{i.i.d.}}$ as $g_{\text{eff}}\to 1^{+}$. As $g_{\text{eff}}$ increases, $K^{\phi}(g_{\text{eff}})$ increases monotonically to $\sim$$1.53\,\text{PR}^{\phi}_{\text{i.i.d.}}$ (Fig. 10a, inset). Thus, for networks with low effective rank, $\text{PR}^{\phi}\sim\alpha\text{PR}^{D}\text{PR}^{\phi}_{\text{i.i.d.}}$ up to an order-one fudge factor that becomes exactly one as $g_{\text{eff}}\to 1^{+}$.
Finally, we consider the dependence of the full spectrum eigenvalues $\lambda^{a}_{i}$ of $\bm{C}^{a}(0)$ on the spectrum of component strengths $D_{a}$. The latter leads to the former via a highly nonlinear relationship. Indeed, sticking to the step-function component-strength spectrum, the ranked eigenvalues $\lambda^{\phi}$ and $\lambda^{x}$ of the covariance matrix decay smoothly rather than exhibiting a cutoff (Fig. 10b). Nevertheless, certain properties of these connectivity and activity spectra are linked. In particular, in the limit $\alpha\text{PR}^{D}\to 0^{+}$, $\alpha\text{PR}^{D}$ and $\text{PR}^{a}$ exhibit a linear relationship (with proportionality factor $K^{a}(g_{\text{eff}})$), implying that the ratios of the second-squared and fourth moments of these spectra are linked (despite that, even in this limit, the full spectra do not coincide).
@@ VI Single-neuron heterogeneity
Figure 5: Effect of single-neuron heterogeneity on dimension of activity in the random-mode model. (a) Dimension of normalized activity $\text{PR}^{\phi}$ versus participation ratio of gain distribution $\text{PR}^{G}$ for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$. Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions. (b) Fractional reduction in dimension for weighted activations relative to $\text{PR}^{\phi}$. Blue: $G_{i}\phi_{i}$; purple: $G^{\text{readout}}_{i}\phi_{i}$. $g_{\text{eff}}=10$. Thick dots: means over 10 simulations, averaging before taking the ratio; lines: theoretical predictions. All simulations use $N=5000$ neurons.
Cortical neurons display broad distributions of firing rates [57]. This heterogeneity among neurons is analogous to the heterogeneity of component strengths in the random-mode model, but operates in the neuron basis rather than the component basis. We now compare how low-dimensional connectivity structure and heterogeneous neuronal properties affect collective activity by extending our framework to include single-neuron heterogeneity.
@@
The single-site picture that determines the two-point functions is largely similar to Eq. 36c.
We analyze collective activity (four-point functions) for (1) unnormalized activations $\Phi_{\bm{\theta}_{i}}(t)=G_{i}\phi(x_{i}(t))$ and (2) normalized activations $\phi_{i}(t)=\phi(x_{i}(t))$, the latter being analogous to $z$-scored firing rates in neural recordings, where $z$-scoring removes single-neuron heterogeneity while preserving structure from cross-neuron and temporal correlations. The four-point functions for the unnormalized and normalized variables are found to be
@@
respectively, where $C^{\Phi}_{12}=q_{2}^{2}C^{\phi}_{12}$ and we have updated the definition of the effective coupling strength to $g_{\text{eff}}=\sqrt{\alpha r_{2}q_{2}}$. Eq. 46b demonstrates that reductions in effective rank, described by $\alpha\text{PR}^{D}$, and reductions in the heterogeneity of gains, described by $\text{PR}^{G}$, have symmetric effects on the dimension of normalized activity, with their combined effect captured by the harmonic mean $\left(\frac{1}{\alpha\text{PR}^{D}}+\frac{1}{\text{PR}^{G}}\right)^{-1}$  [^6] This symmetry arises because the dynamics of the network are determined by an effective coupling matrix $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$, where $\bm{D}=\text{diag}(D_{1},\ldots,D_{M})$ and $\bm{G}=\text{diag}(G_{1},\ldots,G_{N})$. Swapping the distributions of $D_{a}$ and $G_{i}$ does not change the spectrum when $\bm{L}$ and $\bm{R}$ are i.i.d., since $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$ and $\bm{R}^{T}\bm{G}\bm{L}\bm{D}$ have the same nonzero eigenvalues.. One might expect that normalizing the activations before computing the dimension of activity would remove the effect of heterogeneous gains on the dimension. This is not the case, however, since neuronal gains affect the recurrent dynamics, not merely the observed firing rates.
To further explore these effects, we compare the activity dimensions for three sets of activations: unnormalized activations and normalized activations (1 and 2 above); and (3) normalized activations multiplied by an independent set of “readout” gains, $\Phi^{\text{readout}}_{i}(t)=G^{\text{readout}}_{i}\phi_{i}(t)$, where the readout gains have the same distribution as the actual gains. Case (3), whose four-point function is given by Eq. 164, considers variables with the same distribution of firing rates as the unnormalized activations, but where this heterogeneity is unrelated to the recurrent dynamics.
To validate our theoretical predictions (Eq. 46a, Eq. 46b, and Eq. 164), we simulated networks with component strengths and gains given by $D_{a}\propto\exp(-\beta_{D}a/M)$ and $G_{i}\propto\exp(-\beta_{G}i/N)$ (Fig. 5). The dimension of activity is determined by the coupling strength $g_{\text{eff}}$, the effective rank $\alpha\text{PR}^{D}$, and the participation ratio of the gain distribution $\text{PR}^{G}$. Increases in each of these parameters lead to higher activity dimension for all three sets of activations.
@@
In experimental data analysis, it is common to normalize single-neuron activities to control for differences in firing rates across neurons. Our analysis implies that such normalization may not eliminate the effects of single-neuron heterogeneity on collective-activity properties. The persistence of these effects through normalization provides an experimental signature for distinguishing between heterogeneity that affects recurrent dynamics and heterogeneity that affects only the observed firing rates (e.g., in a circuit that is not recurrent at all).

### v3

@@ IV.3 Solution of the random-mode model
where Eq. (37b) indicates that $\eta^{x}(t)$ is a Gaussian field with zero mean and correlation function $\alpha r_{2}C^{\phi}_{\star}(t,t^{\prime})$. These equations are equivalent to those for an i.i.d. model [Eqs. (70)–(72)] with coupling strength
@@
The four-point function, our main object of interest, is (Appendix I.2)
@@
where we have simplified notation by suppressing frequency variables via the shorthand $C^{\phi}_{12}=C^{\phi}_{\star}(\omega_{1})C^{\phi}_{\star}(\omega_{2})$ and $S^{\phi}_{12}=S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})$. The same solution is obtained through a more involved, but intuitive, two-site cavity calculation (Appendix L; Fig. 11).
This solution for the random-mode model is one of our main results, which we now spend some time interpreting.
@@ IV.4 Interpreting the solution
Figure 4: Dimension of activity and collective timescales in the random-mode model. (a) Activity dimension $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ for various coupling strengths $g_{\text{eff}}$. Thin dots, individual simulations; thick dots, means over ten simulations; lines, theoretical predictions. Inset: extended $g_{\text{eff}}=6$ case, showing convergence to i.i.d. coupling behavior with growing $\alpha$. (b) Normalized four-point function $\overline{\Psi^{\phi}(\tau,0)}\equiv\Psi^{\phi}(\tau,0)/\Psi^{\phi}(0,0)$ for various $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$. Inset: theory curves for all $\alpha\text{PR}^{D}$, demonstrating relative invariance of collective timescales to the effective rank. Simulations use $N=5000$ neurons.
@@
Our analysis reveals two features of connectivity that characterize how it shapes activity, namely, the effective coupling strength $g_{\text{eff}}$ and the effective rank $\alpha\text{PR}^{D}$. The former characterizes local structure (the magnitudes of individual couplings), while the latter characterizes global structure (the number of large connectivity components). The above equations determining $C_{\star}^{\phi}(\tau)$ and $\Psi^{\phi}_{\star}(\bm{\tau})$ immediately reveal that global structure in the connectivity is invisible at the single-neuron level but shapes collective activity. Specifically, the correlation function $C^{\phi}_{\star}(\tau)$ is fully determined by $g_{\text{eff}}$ and is insensitive to the effective rank of connectivity. This has important implications for neural data analysis. In particular, analyses of single-neuron activity properties may miss signatures of structured connectivity that become apparent only when analyzing collective activity properties among simultaneously recorded neurons.
@@
What is the nature of the dependence of collective-activity properties, such as the dimension of activity, on the effective rank? To plot this relationship and validate the theory, we simulated networks and computed the participation ratio from the empirical equal-time covariance matrix of the activations (see Appendix M for numerical details). We used a component-strength spectrum $D_{a}\propto\exp(-\beta_{D}a/M)$, which yields $\text{PR}^{D}=\beta^{-1}_{D}\tanh(\beta_{D})$. To attain a desired effective rank $\alpha\text{PR}^{D}<1$, we set $\alpha=M/N=1$ and solved for $\beta$; for $\alpha\text{PR}^{D}>1$, we fixed $\text{PR}^{D}=1$ and increased $\alpha$. We adjusted $g_{\text{eff}}$ by rescaling $D_{a}$ uniformly. We find excellent theory-simulation agreement (Fig. 4). Figure 4(a) illustrates how the dimension of activity, $\text{PR}^{\phi}$, varies with $\alpha\text{PR}^{D}$ for different values of $g_{\text{eff}}$. Consistent with i.i.d.-coupling networks, $\text{PR}^{\phi}$ increases monotonically with $g_{\text{eff}}$ [39]. It also increases monotonically with $\alpha\text{PR}^{D}$.
@@
Taking $\alpha\text{PR}^{D}\to\infty$ for fixed $g_{\text{eff}}$ in Eq. (39) recovers the formula for i.i.d. couplings [Eq. (73)] of Ref. [39]. Indeed, in this limit, $\bm{J}$ approaches an i.i.d. matrix. Correspondingly, $\text{PR}^{\phi}$ approaches the activity dimension of a network with i.i.d. couplings [Fig. 4(a) inset]. Comparing Eq. (73) to the corresponding i.i.d. formula ($1/\alpha\text{PR}^{D}=0$) and noting that single-neuron properties are unaffected by $\alpha\text{PR}^{D}$, one sees that any structure beyond i.i.d. connectivity ($1/\alpha\text{PR}^{D}$ finite) strictly decreases the dimension of activity.
@@
The dimension of activity depends on $\Psi^{\phi}_{\star}(\bm{\tau})$ at $\bm{\tau}=(0,0)$; we now examine the temporal profile of this function. Figure 4(b) shows a normalized version of $\Psi^{\phi}(\tau,0)$, which is related to the correlation functions of leading principal components, for various values of $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$. The decay timescale of $\Psi^{\phi}_{\star}(\tau,0)$ decreases with increasing $g_{\text{eff}}$, approaching a limiting behavior at $g_{\text{eff}}\approx 10$. While $\alpha\text{PR}^{D}$ affects the overall scale of this function and, consequently, the dimension of activity, it has little effect on the decay timescale and, hence, little effect on the timescales of leading principal components (Fig. 4(b), inset). Thus, the effective rank primarily influences the dimension of activity rather than the collective temporal structure.
@@
A remarkable feature of these results is that only two connectivity statistics, $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$, are sufficient to determine the two- and four-point functions. Because of the nonlinear, high-dimensional nature of the network, this reduction is unexpected. This reduction breaks down when structured mode overlaps are introduced (Sec. VII), in which case the formula for the four-point function requires knowledge of the full joint distribution of component strengths and overlaps.
@@ V Limiting behavior of the random-mode model
Here we examine the behavior of the random-mode model in two limiting cases: as the system approaches the transition to chaos from above ($g_{\text{eff}}\to 1^{+}$) and in the limit of low effective rank ($\alpha\text{PR}^{D}\to 0^{+}$).
We first consider the limit $g_{\text{eff}}\to 1^{+}$ for arbitrary $\alpha\text{PR}^{D}$. The relevant expansion parameter is $0<g_{\text{eff}}-1\ll 1$. For networks with i.i.d. couplings with coupling strength $g$, it is known that, as $g\to 1^{+}$, $\Psi^{a}_{\star,\text{i.i.d.}}(\bm{\tau})\sim 1/(g-1)$ and $\text{PR}^{a}_{\text{i.i.d.}}\sim(g-1)^{3}$ [39]. For the random-mode model, to leading order in $1/(g_{\text{eff}}-1)$,
@@
The difference between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$ is a higher-order correction, so we also have $\text{PR}^{a}=\text{PR}^{S}\text{PR}^{a}_{\text{i.i.d.}}$; that is, in the limits of both low effective rank and small effective coupling strength, the dimension of activity is equal to the dimension of the singular-value spectrum multiplied by the dimension of activity for an i.i.d. network.
@@
We verify this limiting behavior in simulations using $\alpha=1$ and a step function for the component strengths: $D_{a}=\text{const.}$ for $a/N\leq\text{PR}^{D}$ and $D_{a}=0$ otherwise. We varied $\text{PR}^{D}$, observing linear relationships with $\text{PR}^{\phi}$ and $\text{PR}^{x}$ in the $\text{PR}^{D}\to 0^{+}$ limit [Fig. 10(a)].
Focusing on $K^{\phi}(g_{\text{eff}})$, note that its expression [Eq. (44)] resembles the formula for the dimension of activity in an i.i.d. network with coupling strength $g_{\text{eff}}$ [Eq. (73)], but includes an additional factor $|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}$ in the integral. Numerical evaluation confirms that, as expected from Eq. (42), $K^{\phi}(g_{\text{eff}})\to\text{PR}^{\phi}_{\text{i.i.d.}}$ as $g_{\text{eff}}\to 1^{+}$. As $g_{\text{eff}}$ increases, $K^{\phi}(g_{\text{eff}})$ increases monotonically to approximately $1.53\,\text{PR}^{\phi}_{\text{i.i.d.}}$ [Fig. 10(a), inset]. Thus, for networks with low effective rank, $\text{PR}^{\phi}\sim\alpha\text{PR}^{D}\text{PR}^{\phi}_{\text{i.i.d.}}$ up to an order-one fudge factor that becomes exactly one as $g_{\text{eff}}\to 1^{+}$.
Finally, we consider the dependence of the full eigenvalue spectrum $\lambda^{a}_{i}$ of $\bm{C}^{a}(0)$ on the spectrum of component strengths $D_{a}$. The latter leads to the former via a highly nonlinear relationship. Indeed, using the same step-function component-strength spectrum, the ranked eigenvalues $\lambda^{\phi}$ and $\lambda^{x}$ of the covariance matrix decay smoothly rather than exhibiting a cutoff [Fig. 10(b)]. Nevertheless, certain properties of these connectivity and activity spectra are linked. In particular, in the limit $\alpha\text{PR}^{D}\to 0^{+}$, $\alpha\text{PR}^{D}$ and $\text{PR}^{a}$ exhibit a linear relationship [with proportionality factor $K^{a}(g_{\text{eff}})$], implying that the ratios of the squared second moments and fourth moments of these spectra are linked (despite, even in this limit, the full spectra not coinciding).
@@ VI Single-neuron heterogeneity
Figure 5: Effect of single-neuron heterogeneity on dimension of activity in the random-mode model. (a) Dimension of normalized activity $\text{PR}^{\phi}$ versus participation ratio of gain distribution $\text{PR}^{G}$ for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$. Thin dots, individual simulations; thick dots, means over ten simulations; lines, theoretical predictions. (b) Fractional reduction in dimension for weighted activations relative to $\text{PR}^{\phi}$. Blue, $G_{i}\phi_{i}$; purple, $G^{\text{readout}}_{i}\phi_{i}$. $g_{\text{eff}}=10$. Thick dots, means over ten simulations, averaging before taking the ratio; lines, theoretical predictions. All simulations use $N=5000$ neurons.
Cortical neurons display broad distributions of firing rates [66]. This heterogeneity among neurons is analogous to the heterogeneity of component strengths in the random-mode model, but operates in the neuron basis rather than the component basis. We now compare how low dimensional connectivity structure and heterogeneous neuronal properties affect collective activity by extending our framework to include single-neuron heterogeneity.
@@
The single-site picture that determines the two-point functions is largely similar to Eq. (37c).
We analyze collective activity (four-point functions) for (1) unnormalized activations $\Phi_{\bm{\theta}_{i}}(t)=G_{i}\phi\bm{(}x_{i}(t)\bm{)}$ and (2) normalized activations $\phi_{i}(t)=\phi\bm{(}x_{i}(t)\bm{)}$, the latter being analogous to $z$-scored firing rates in neural recordings, where $z$ scoring removes single-neuron heterogeneity while preserving structure from cross-neuron and temporal correlations. The four-point functions for the unnormalized and normalized variables are found to be
@@
respectively, where $C^{\Phi}_{12}=q_{2}^{2}C^{\phi}_{12}$ and we have updated the definition of the effective coupling strength to $g_{\text{eff}}=\sqrt{\alpha r_{2}q_{2}}$. Equation (47b) demonstrates that reductions in effective rank, described by $\alpha\text{PR}^{D}$, and reductions in the heterogeneity of gains, described by $\text{PR}^{G}$, have symmetric effects on the dimension of normalized activity, with their combined effect captured by the harmonic mean $\left(\frac{1}{\alpha\text{PR}^{D}}+\frac{1}{\text{PR}^{G}}\right)^{-1}$  [^6] This symmetry arises because the dynamics of the network are determined by an effective coupling matrix $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$, where $\bm{D}=\text{diag}(D_{1},\ldots,D_{M})$ and $\bm{G}=\text{diag}(G_{1},\ldots,G_{N})$. Swapping the distributions of $D_{a}$ and $G_{i}$ does not change the spectrum when $\bm{L}$ and $\bm{R}$ are i.i.d., since $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$ and $\bm{R}^{T}\bm{G}\bm{L}\bm{D}$ have the same nonzero eigenvalues.. One might expect that normalizing the activations before computing the dimension of activity would remove the effect of heterogeneous gains on the dimension. This is not the case, however, since neuronal gains affect the recurrent dynamics, not merely the observed firing rates.
To further explore these effects, we compare the activity dimensions for three sets of activations: unnormalized activations and normalized activations (1 and 2 above); and (3) normalized activations multiplied by an independent set of “readout” gains, $\Phi^{\text{readout}}_{i}(t)=G^{\text{readout}}_{i}\phi_{i}(t)$, where the readout gains have the same distribution as the actual gains. Case (3), whose four-point function is given by Eq. (166), considers variables with the same distribution of firing rates as the unnormalized activations, but where this heterogeneity is unrelated to the recurrent dynamics.
To validate our theoretical predictions [Eqs. (47a), (47b), and (166)], we simulated networks with component strengths and gains given by $D_{a}\propto\exp(-\beta_{D}a/M)$ and $G_{i}\propto\exp(-\beta_{G}i/N)$ (Fig. 5). The dimension of activity is determined by the coupling strength $g_{\text{eff}}$, the effective rank $\alpha\text{PR}^{D}$, and the participation ratio of the gain distribution $\text{PR}^{G}$. Increases in each of these parameters lead to a higher activity dimension for all three sets of activations.
@@
In experimental data analysis, it is common to normalize single-neuron activities to control for differences in firing rates across neurons. Our analysis implies that such normalization may not eliminate the effects of single-neuron heterogeneity on collective-activity properties. The persistence of these effects through normalization provides an experimental signature for distinguishing between heterogeneity that affects recurrent dynamics and heterogeneity that affects only the observed firing rates, for example, in a circuit that is not recurrent.